OpenStudy (danielajohana):

What is the maximum volume of a box that can be made from an 11 inch by 18 inch piece of paper? Show your work Will fan and medal

OpenStudy (campbell_st):

let x he the height of the box... length = 18 - 2x width = 11 - 2x  the volume is \[V = x (18-2x)(11-2x)\] multiply it out